OpenStudy (anonymous):

jose found a board to make a shelf. it was 4 1/7 ft long. he needed 1/3 of it for his shelf. how long was the shelf??

OpenStudy (anonymous):

4 +1/7 = 29/7 29/7*1/3 = 29/21 = 1.38